# Topologically noise robust network steering without inputs

### v1

@@ II.2 Topologically Robust Swap Steering
$p(\{a\}_{n-1/1}|\{\lambda\},a_{1})=p(\{a\}_{n-1/1}|\lambda_{1},\lambda_{n})$ (23)
@@
Consequently, the observed swap-steering inequality from the violation of (II.2) is robust against any network structure among the untrusted parties. This would also include cases when the untrusted parties might communicate with each other. Consequently, one can detect a quantum advantage in a network against much more powerful classical models, which is impossible in the standard multipartite quantum steering scenarios. Also, notice from fact 3 that the gap between the quantum and SOHS values is considerably large and thus it can be observed in experiments. Consequently, this is the first example of a topologically noise robust advantage in a quantum network.
II.3 All bipartite entangled states generate swap-steering correlations in the $n-$ring network
@@
where $p(0,0,\ldots,0|st)$ is the probability of obtaining outcome 00 by all the parties given the input $s,t$ of the trusted party $A_{1}$. Let us now find its SOHS bound.
@@
The proof of the above fact is in the Appendix. Let us now consider that the source connecting the parties $A_{2},A_{3}$ generates the state $\tilde{\rho}_{A_{2}^{1},A_{3}^{2}}\in\mathbb{C}^{{d_{A_{1}^{1}}}}\otimes\mathbb{C}^{d_{A_{1}^{2}}}$. Now the source connecting the trusted party $A_{1}$ with $A_{2}$ produces the maximally entangled state of dimension $d_{A_{1}^{1}}$, that is, $|\psi_{A_{1}^{1},A_{2}^{2}}\rangle=|\phi^{+}_{d_{A_{1}^{1}}}\rangle$. Rest of the other sources generate the maximally entangled state $|\phi^{+}_{d_{A_{1}^{2}}}\rangle$. Moreover, $A_{2}$ performs the measurements $\{|\phi^{+}\rangle\!\!\langle\phi^{+}|_{d_{A_{1}^{1}}},\mathbbm{1}-|\phi^{+}\rangle\!\!\langle\phi^{+}|_{d_{A_{1}^{1}}}\},$ and rest of the parties $A_{3},\ldots,A_{n}$ perform $\{|\phi^{+}\rangle\!\!\langle\phi^{+}|_{d_{A_{1}^{2}}},\mathbbm{1}-|\phi^{+}\rangle\!\!\langle\phi^{+}|_{d_{A_{1}^{2}}}\}$ respectively. Now, consider a party $A_{i}$ for $i=4,\ldots,n$ obtains outcome 00, then the post-measurement state between $A_{i-1},A_{i+1}(n+1\equiv 1)$ is the maximally state of dimension $d_{A_{1}^{2}}$. Consequently, it is straightforward to observe that the $n-$network reduces to a triangle network such that $A_{1},A_{3}$ share $|\phi^{+}_{d_{A_{1}^{2}}}\rangle$. Now, $A_{2},A_{3}$ projecting the joint state over their $0-$th outcome swaps the state $\tilde{\rho}$ to the trusted $A_{1}$ and thus one is left with the expression
@@
Notice that the factor $1/{d_{A_{1}^{1}}^{2}d_{A_{1}^{2}}^{n-1}}$ takes into account the probability of obtaining outcome 00 by all the parties. Consequently, every entangled state results in $n-$ring swap-steerable correlations. It is important to note here that although the above prescription closely follows [48], the sources connected to the trusted party are also trusted in [48] which is not the case here.
III Discussions
@@
Acknowledgements—We thank Stefano Pironio for mentoring this work and providing crucial insights. This project was funded within the QuantERA II Programme (VERIqTAS project) that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733.
@@ Appendix A Proofs in the Triangle scenario
[…]
@@ A.2 Triangle network weak self-testing
Assume that Alice, Bob and Charlie observe that the correlations $p(a,b,c)$ in the steering functional $W_{3}$ are equal (I.2) along with $W_{3}=1$ such that trusted Alice performs the Bell-basis. Then, the following statements hold true:
@@ 1.    There exist unitary transformation $U$, such that the state generated by the sources are certified as
$(\mathbbm{1}_{A_{1}}\otimes U_{B_{2}})|\psi_{A_{1}B_{2}}\rangle=(\mathbbm{1}_{A_{2}}\otimes U_{C_{1}})|\psi_{A_{2}C_{1}}\rangle=(\mathbbm{1}_{B_{1}}\otimes U_{C_{2}})|\psi_{B_{1}C_{2}}\rangle=|\phi^{+}\rangle.$ (30)
@@ 2.    The measurement of Bob $\{M_{b}\}$ and Charlie $\{M_{c}\}$ is certified as
$1\displaystyle 4\sum_{b}p(0,b,b)=\frac{1}{4}\sum_{x=0}^{3}\langle\phi^{+}_{4}|B_{0}^{x}\otimes C_{0}^{4-x}|\phi^{+}_{4}\rangle=$ 1 (45)
@@
As all the operators inside the <.><.>< . > are unitary, we have that
@@
Thus, $(\mathbbm{1}\otimes\sigma_{z})|\xi\rangle$ is an eigenvector of $B_{0}$ with eigenvalue $-1$ where $\sigma_{z}=|0\rangle\!\!\langle 0|+|1\rangle\!\!\langle 1|$.
@@
Thus, $(\mathbbm{1}\otimes\sigma_{z}\sigma_{x})|\xi\rangle$ is an eigenvector of $B_{0}$ with eigenvalue $-i$.
@@ Appendix B Proofs for the $n-$ring scenario
B.1 Local bound of n-party swap steering witness
@@
[…]
@@
[…]
@@ B.3 Every bipartite entangled state is swap-steerable
where $p(0,0,\ldots,0|st)$ is the probability of obtaining outcome 00 by all the parties given the input $s,t$ of the trusted party $A_{1}$.
@@
$\mathcal{S}_{\tilde{\rho}}$ $=$ $\sum_{\{\lambda\}}\prod_{j=1}^{n}p(\lambda_{j})\Gamma(\rho_{\lambda_{1}}\otimes\rho_{\lambda_{n}})\prod_{i=2}^{n}p(0,\ldots,0|\{\lambda_{i}\})$ (68)

### v2

@@ III.1 Topologically Robust Swap Steering
$p(\{a\}_{n-1/1}|\{\lambda\},a_{1})=p(\{a\}_{n-1/1}|\lambda_{1},\lambda_{n})\ \$(23)
@@
Consequently, the observed swap-steering inequality from the violation of (III.1) is robust against any network structure among the untrusted parties. This would also include cases when the untrusted parties might communicate with each other. Consequently, one can detect a quantum advantage in a network against much more powerful classical models, which is impossible in the standard multipartite quantum steering scenarios. Also, notice from Fact 3 that the gap between the quantum and SOHS values is considerably large and thus it can be observed in experiments. Consequently, this is the first example of a topologically noise robust advantage in a quantum network.
Notice that with no information about the network structure, observing correlations compatible with $n$ party steering does not imply that the correlation was generated genuinely by $n$ different parties. This is simply an inverse of the topological robustness property of swap-steering. For example, consider the experiment on the minimal scenario [17] with a single untrusted party generating maximal swap-steering violation. The untrusted party can simply use the refinement scheme [Eq. (III)] used to extend the witness to the $n$ party scenario to his output and report a n1n-1 string bit. This by definition, will result in statistics that maximally violate the $n$ party witness $W_{n}$. Thus, one cannot naively conclude sucessful $n$ party steering by only observing the correct statistics, it must be supplemented with some information about the network structure. Note that this is a core feature of topological robustness which allows us to observe steering without any information about the network structure among untrusted parties.
III.2 All bipartite entangled states generate swap-steering correlations in the $n$-ring network
@@
where $p(0,0,\ldots,0|st)$ is the probability of obtaining outcome 0 by all the parties given the input $s,t$ of the trusted party $A_{1}$. Let us now find its SOHS bound.
@@
The proof of the above fact is in the Appendix. Let us now consider that the source connecting the parties $A_{2},A_{3}$ generates the state $\tilde{\rho}_{A_{2}^{1},A_{3}^{2}}\in\mathbb{C}^{{d_{A_{1}^{1}}}}\otimes\mathbb{C}^{d_{A_{1}^{2}}}$. Now the source connecting the trusted party $A_{1}$ with $A_{2}$ produces the maximally entangled state of dimension $d_{A_{1}^{1}}$, that is, $|\psi_{A_{1}^{1},A_{2}^{2}}\rangle=|\phi^{+}_{d_{A_{1}^{1}}}\rangle$. Rest of the other sources generate the maximally entangled state $|\phi^{+}_{d_{A_{1}^{2}}}\rangle$. Moreover, $A_{2}$ performs the measurements $\{|\phi^{+}\rangle\!\!\langle\phi^{+}|_{d_{A_{1}^{1}}},\mathbbm{1}-|\phi^{+}\rangle\!\!\langle\phi^{+}|_{d_{A_{1}^{1}}}\},$ and rest of the parties $A_{3},\ldots,A_{n}$ perform $\{|\phi^{+}\rangle\!\!\langle\phi^{+}|_{d_{A_{1}^{2}}},\mathbbm{1}-|\phi^{+}\rangle\!\!\langle\phi^{+}|_{d_{A_{1}^{2}}}\}$ respectively. Now, consider a party $A_{i}$ for $i=4,\ldots,n$ obtains outcome 0, then the post-measurement state between $A_{i-1},A_{i+1}(n+1\equiv 1)$ is the maximally state of dimension $d_{A_{1}^{2}}$. Consequently, it is straightforward to observe that the $n$-network reduces to a triangle network such that $A_{1},A_{3}$ share $|\phi^{+}_{d_{A_{1}^{2}}}\rangle$. Now, $A_{2},A_{3}$ projecting the joint state over their 00-th outcome swaps the state $\tilde{\rho}$ to the trusted $A_{1}$ and thus one is left with the expression
@@
Notice that the factor $1/{d_{A_{1}^{1}}^{2}d_{A_{1}^{2}}^{n-1}}$ takes into account the probability of obtaining outcome 0 by all the parties. Consequently, every entangled state results in $n$-ring swap-steerable correlations. It is important to note here that although the above prescription closely follows [48], the sources connected to the trusted party are also trusted in [48] which is not the case here.
IV Discussions
@@
V Acknowledgements
We thank Stefano Pironio for mentoring this work and providing crucial insights. This project was funded within the QuantERA II Programme (VERIqTAS project) that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733. S.S. also acknowledges National Science Centre, Poland, grant Opus 25, 2023/49/B/ST2/02468.
@@ Appendix A Proofs in the Triangle scenario
$c_{a_{1},a_{2},a_{3}}=\begin{cases}1\ \mathrm{if}\ a_{1}-a_{2}-(-1)^{a_{2}}a_{3}\ \mathrm{mod}\ 4=0\\ 0\ \mathrm{otherwise}.\end{cases}$ (26)
@@ A.2 Triangle network weak self-testing
Assume that Alice, Bob and Charlie observe that the correlations $p(a,b,c)$ in the steering functional $W_{3}$ are equal [see Eq. (60) below for $n=3$] along with $W_{3}=1$ such that trusted Alice performs the Bell-basis. Considering identical sources, such that the states generated by them are pure and measurements of untrusted parties are projective, the following statements hold true:
@@ 1. There exist unitary transformation $U$, such that the state generated by the sources are certified as
$(\mathbbm{1}_{A_{1}}\otimes U_{B_{2}})|\psi_{A_{1}B_{2}}\rangle=(\mathbbm{1}_{A_{2}}\otimes U_{C_{1}})|\psi_{A_{2}C_{1}}\rangle=(\mathbbm{1}_{B_{1}}\otimes U_{C_{2}})|\psi_{B_{1}C_{2}}\rangle=|\phi^{+}\rangle.\ \$(30)
@@ 2. The measurement of Bob $\{M_{b}\}$ and Charlie $\{M_{c}\}$ is certified as
$4\sum_{b}p(0,b,b)=\frac{1}{4}\sum_{x=0}^{3}\langle\phi^{+}_{4}|B_{0}^{x}\otimes C_{0}^{4-x}|\phi^{+}_{4}\rangle=1$ (45)
@@
As all the operators inside the <.><.> are unitary, we have that
@@
Thus, $(\mathbbm{1}\otimes\sigma_{z})|\xi\rangle$ is an eigenvector of $B_{0}$ with eigenvalue 1-1 where $\sigma_{z}=|0\rangle\!\!\langle 0|+|1\rangle\!\!\langle 1|$.
@@
Thus, $(\mathbbm{1}\otimes\sigma_{z}\sigma_{x})|\xi\rangle$ is an eigenvector of $B_{0}$ with eigenvalue i-i.
@@ Appendix B Proofs for the $n$-ring scenario
B.1 SOHS and quantum bounds of $n$-party swap steering witness
@@
$c_{\{a\}_{n}}=\begin{cases}1\ \mathrm{if}\ a_{1}-a_{2}-\sum_{i=3}^{n}(-1)^{\sum_{k=2}^{i-1}a_{k}}a_{i}\ \mathrm{mod}\ 4=0\\ 0\ \mathrm{otherwise}.\end{cases}$ (60)
@@
From the condition (60), we have that the terms in the witness $W_{n}$ must satisfy $a_{1}-a_{2}-\sum_{i=3}^{n}(-1)^{\sum_{k=2}^{i-1}a_{k}}a_{i}\ \mathrm{mod}\ 4=0$. Observe that for every $\{a\}_{n-1/1}$, there exists a unique solution $a_{1}$, which implies that all combinations of n1n-1 strings $\{a\}_{n-1/1}$ appear exactly once in the summation. Thus, for the topologically robust version of the SOHS model (III.1), the witness $W_{n}$ can be simplified as
@@
Fact 2.
The maximal quantum value of the witness $W_{n}$ in the $n$-ring network with one trusted party is $1$.
Proof.
For this proof we will use the following notation $|\phi_{+}\rangle\rightarrow|0\rangle$, $|\psi_{+}\rangle\rightarrow|1\rangle$, $|\phi_{-}\rangle\rightarrow|2\rangle$ and $|\psi_{-}\rangle\rightarrow|3\rangle$. The $n$ party state is then written as
$|\psi_{n}\rangle=\prod_{i}^{n}|0\rangle_{A_{i}^{2},A_{i+1}^{1}}$ (63)
One can easily check the following properties of bell states given below.
$|0\rangle_{1,2}|0\rangle_{3,4}=\frac{\left(|0\rangle|0\rangle+|1\rangle|1\rangle+|2\rangle|2\rangle+|3\rangle|3\rangle\right)_{1,4,2,3}}{2}$
$|0\rangle_{1,2}|1\rangle_{3,4}=\frac{\left(|1\rangle|0\rangle+|0\rangle|1\rangle+|2\rangle|3\rangle+|3\rangle|2\rangle\right)_{1,4,2,3}}{2}$
$|0\rangle_{1,2}|2\rangle_{3,4}=\frac{\left(|2\rangle|0\rangle+|0\rangle|2\rangle-|1\rangle|3\rangle-|3\rangle|1\rangle\right)_{1,4,2,3}}{2}$
$|0\rangle_{1,2}|3\rangle_{3,4}=\frac{\left(|3\rangle|0\rangle-|0\rangle|3\rangle+|1\rangle|2\rangle-|2\rangle|1\rangle\right)_{1,4,2,3}}{2}$ (64)
These equations are the key component of the proof. Notice how the state of the second subsystem determines the form of the permuted state in the same manner as the reverse coarse-graining scheme used to construct the witness. We will now prove that $p(\{a\}_{n})=c_{\{a\}_{n}}/4^{n-1}$ for the $n$ partite state (63). Using (B.1) we have,
$|\psi_{n}\rangle=\left(\prod_{i}^{n-2}|0\rangle_{A_{i}^{2},A_{i+1}^{1}}\right)|0\rangle_{A_{n-1}^{2},A_{n}^{1}}|0\rangle_{A_{n}^{2},A_{1}^{1}}=\left(\prod_{i}^{n-2}|0\rangle_{A_{i}^{2},A_{i+1}^{1}}\right)\frac{1}{2}\left(|0\rangle|0\rangle+|1\rangle|1\rangle+|2\rangle|2\rangle+|3\rangle|3\rangle\right)_{A_{n-1}^{2},A_{1}^{1},A_{n}^{1},A_{n}^{2}}.\quad$ (65)
Suppose $n=2$. The product on the left reduces to identity and the sum of states on the right produce the desired correlation $p(\{a\}_{2})=c_{\{a\}_{2}}/4$. Repeating this step, we have
$|\psi_{n}\rangle=\left(\prod_{i}^{n-3}|0\rangle_{A_{i}^{2},A_{i+1}^{1}}\right)\frac{1}{4}\left(|0\rangle|0\rangle|0\rangle+|0\rangle|1\rangle|1\rangle+|0\rangle|2\rangle|2\rangle+|0\rangle|3\rangle|3\rangle+|1\rangle|1\rangle|0\rangle+|1\rangle|0\rangle|1\rangle+|1\rangle|2\rangle|3\rangle+|1\rangle|3\rangle|2\rangle+\right$.
$\left|2\rangle|2\rangle|0\rangle+|2\rangle|0\rangle|2\rangle-|2\rangle|1\rangle|3\rangle-|2\rangle|3\rangle|1\rangle+|3\rangle|3\rangle|0\rangle-|3\rangle|0\rangle|3\rangle+|3\rangle|1\rangle|2\rangle-|3\rangle|2\rangle|1\rangle\right)_{A_{n-2}^{2},A_{1}^{1},A_{n-1}^{1},A_{n-1}^{2},A_{n}^{1},A_{n}^{2}}.\qquad$ (66)
Again, suppose that $n=3$. For the case of the triangle network, the product of the states on the left reduces to identity and the sum of states on the right reproduces the desired behavior $p(\{a\}_{3})=c_{\{a\}_{3}}/16$. Further applications of the basis swap on the distributed state will respect the reverse coarse-graining scheme, which by definition, will produce terms of the n-ring witness. Each term of the n-ring witness will appear in the sum with equal probability $1/4^{n-1}$ (easily checked for the 2 and 3-party witness). This concludes the proof. ∎
@@ B.3 Every bipartite entangled state is swap-steerable
where $p(0,0,\ldots,0|st)$ is the probability of obtaining outcome 0 by all the parties given the input $s,t$ of the trusted party $A_{1}$.
@@
$\mathcal{S}_{\tilde{\rho}}$ $=$ $\sum_{\{\lambda\}}\prod_{j=1}^{n}p(\lambda_{j})\Gamma(\rho_{\lambda_{1}}\otimes\rho_{\lambda_{n}})\prod_{i=2}^{n}p(0,\ldots,0|\{\lambda_{i}\})\ \$(72)